# Holographic Screen Sequestration

### 11 Dec 2023 v1

@@ I Introduction
The past quarter-century of the anti-de Sitter/conformal field theory (AdS/CFT) correspondence Maldacena:1997re ; Gubser:1998bc ; Witten:1998qj ; Aharony:1999ti has ushered in a revolution of our understanting of gravitation and spacetime. Rather than being fundamental concepts, in AdS/CFT both quantum gravity and the bulk dimension of spacetime itself can be seen as emergent phenomena Susskind:1994vu ; Bousso:2002ju ; Faulkner:2013ica , arising as effective descriptions encoded via complicated and subtle details of the dynamics in the nongravitational world of the boundary. This encoding has been understood as a form of quantum error correction Almheiri:2014lwa ; Pastawski:2015qua , with the geometry encoding entanglement—specifically, the fine grained (von Neumann) entropy—through areas of surfaces, or more general quantities in a gravitational effective field theory Dong:2013qoa . [^1] See Refs. Dong:2023bax ; Gesteau:2023hbq for recent analyses of the relationship between renormalization and holographic entropy in effective theories. The Ryu-Takayanagi (RT) formula Ryu:2006bv , and more generally the Hubeny-Rangamani-Takayanagi (HRT) formalism Hubeny:2007xt for dynamical spacetimes, give the prescriptions for finding the appropriate bulk surfaces whose areas compute the entropy corresponding to a given boundary region, generalizing the celebrated Bekenstein-Hawking entropy of black holes Bekenstein:1972tm ; Bekenstein:1973ur ; Hawking:1976de .
@@
Despite the success of these famous results and many others, significant puzzles remain. How does spacetime emerge when it is not anti-de Sitter? That is, what is the analogue of the boundary theory for general spacetimes, and where is it located? As a starting point, one can look to history, where Hawking’s area theorem for black holes arguably was a critical seed leading eventually to our current holographic understanding of AdS/CFT. Happily, new area theorems for general spacetimes have been found by Bousso and Engelhardt Bousso:2015mqa ; Bousso:2015qqa , in the form of monotonic area increase along a special hypersurface called a holographic screen. The holographic screen has been conjectured to be the appropriate analogue of the AdS boundary for a holographic description of general spacetimes Bousso:1999cb ; Bousso:2015mqa ; Bousso:2015qqa ; Nomura:2017npr . Within AdS/CFT, an entropic interpretation of the outermost spacelike portion of a holographic screen, that is, an apparent horizon, was provided by Refs. Engelhardt:2017aux ; Engelhardt:2018kcs . There it was shown that the area of a slice of the screen is equal to the outer entropy: the area of the maximal HRT surface, in units of $4G\hbar$, subject to holding the geometry in the outer wedge—the exterior of the slice of the screen—fixed and marginalizing over all possible completions of the spacetime. In this sense, at least the apparent horizon can be viewed as a coarse grained holographic measure of entanglement entropy. Moreover, the outer entropy can be viewed as computing a quasilocal energy for general (i.e., nonmarginal) surfaces Bousso:2018fou ; Wang:2020vxc .
@@
It is this latter question that we answer in this paper. Specifically, we show that holographic screens are sequestered relative to the HRT surface: given an HRT surface, a holographic screen is contained entirely within either its interior, exterior, past, or future, being forever forbidden from crossing the null congruences launched from the HRT surface. This paper is organized as follows. After providing a self-contained primer of the geometric background and definitions in Sec. II, we prove our main sequestration result, Theorem III.6, in Sec. III. There are special cases in which, under certain conditions, the screen can touch—but not pass through—a null congruence from the HRT surface, and we enumerate the possibilities in detail in Sec. III.3. We comment on obstructions to straightforward application of a coarse grained interpretation of timelike holographic screens in Sec. IV. We conclude in Sec. V, placing our result in context and commenting on connections to non-HRT extremal surfaces (e.g., “minimax” surfaces) and semiclassical quantum corrections.
@@ II.1 Topological definitions
Throughout this paper, we will be concerned with geometric objects constructed from light rays: causally-defined regions, expansion of bundles of light rays, etc. We begin, for completeness, by reviewing some terminology, following Refs. Engelhardt:2018kcs ; Nomura:2018aus ; Bousso:2018fou . For a set of points $S\subset{\cal M}$, we define the chronological and causal future of $S$ as $I^{+}[S]$ and $J^{+}[S]$, denoting the sets of all points in ${\cal M}$ connected with any point in $S$ by timelike or null paths, respectively. The chronological and causal pasts $I^{-}[S]$ and $J^{-}[S]$ are defined analogously. We will assume that ${\cal M}$ is globally hyperbolic: it is free of closed causal curves and, for all points $p,q\in{\cal M}$, $J^{+}(p)\cap J^{-}(q)$ is compact (or, in the asymptotically-AdS case, when supplemented with boundary conditions Avis:1977yn ). We define the future (past) domain of dependence of $S$, $D^{\pm}[S]$, as the set of all $p\in{\cal M}$ for which all past (future) inextendible causal curves through $p$ pass through $S$, and define $D[S]=D^{+}[S]\cup D^{-}[S]$. We will refer to a surface as acausal if every pair of points on the surface is spacelike-separated; the weaker condition achronal is defined analogously, but allowing null-separated points. A Cauchy slice $\Sigma$ is defined as an achronal codimension-one hypersurface for which $D[\Sigma]={\cal M}$. We will at times refer to an achronal slice of the boundary ${\cal B}=\partial{\cal M}$ (i.e., ${\cal B}\cap\Sigma=\partial\Sigma$ for a Cauchy slice $\Sigma$) as itself a boundary $B$, trusting that the difference between $B$ and ${\cal B}$ will be clear in context. From a closed, codimension-two, achronal surface $\nu\subset{\cal M}$, we can define four orthogonal null congruences—families of null geodesics hereafter referred to as light sheets—associated with the null rays launched either outward or inward, past- or future-directed, from $\nu$, spatially orthogonal to $\nu$ at the launch point. Let us define the null vector on one side (e.g., “outward”) as $k$ and on the other side (“inward”) as $\ell$, with $+k$ and $+\ell$ being future-directed, and write the four congruences as $N_{\pm k}[\nu]$ and $N_{\pm\ell}[\nu]$. When convenient, we will also denote the union of future- and past-directed null congruences by $N_{k}[\nu]=N_{+k}[\nu]\cup N_{-k}[\nu]$ and $N_{\ell}[\nu]=N_{+\ell}[\nu]\cup N_{-\ell}[\nu]$. Geodesics can exit the light sheets if Wald ; Penrose and only if Akers:2017nrr they encounter either a caustic or an intersection with a distinct null geodesic. For $\nu$ Cauchy-splitting—that is, dividing a Cauchy surface $\Sigma\supset\nu$ into two sides $\Sigma^{\pm}$—we can write
@@
See Fig. 1. We will choose $\Sigma^{+}[\nu]$ to be on the side of $\nu$ toward which $k$ points and $\Sigma^{-}[\nu]$ to be on the side toward which $\ell$ points. We define the outer wedge $O_{W}[\nu]=D[\Sigma^{+}[\nu]]$ and inner wedge $I_{W}[\nu]=D[\Sigma^{-}[\nu]]$.
@@ II.2 Dynamical definitions
Having concluded the topological preliminaries, we now turn to dynamics. We will define the 2+2 formalism for general relativity, in which the Einstein equations governing the evolution of spacetime can be recast in terms of these light-centric objects. This formulation is known as the characteristic initial data formalism for general relativity Rendall:2000 ; Brady:1995na ; ChoquetBruhat:2010ih ; Luk:2011vf ; Chrusciel:2012ap ; Chrusciel:2012xf ; Chrusciel:2014lha , and one can view it as akin to the more familiar 3+1 Arnowitt-Deser-Misner formalism, but with spatial Cauchy data replaced with Cauchy data on a null surface. Given a Cauchy-splitting surface $\nu$, let us focus on the light sheet along the $k$-direction, $N_{k}[\nu]$. We will define $k$ to be affinely parameterized as it is parallel transported along itself, and $\ell$ to be parallel transported along $k$ but continually rescaled so that $k\cdot\ell=-1$. We define the induced metric as $q_{\mu\nu}=g_{\mu\nu}+k_{\mu}\ell_{\nu}+k_{\nu}\ell_{\mu}$, in terms of which we have the null extrinsic curvature,
@@ II.3 Screen-related definitions
We define a Hubeny-Rangamani-Takayanagi (HRT) surface as an extremal surface $X$ homologous to the boundary $B$, such that there exists a Cauchy slice $\Sigma$ of $D[O_{W}[X]]$ on which $X$ is minimal. As such surfaces can be identified via a maximin optimization Wall:2012uf , we will sometimes refer to HRT surfaces as “being maximin.” Though HRT surfaces can be defined for subregions of the boundary, throughout, we will be interested in taking $B$ to be a complete connected component. One can show that HRT surfaces are minimar Engelhardt:2018kcs . HRT surfaces are of particular interest to holography, since the area of the HRT suface, divided by $4G\hbar$, corresponds to the fine grained (von Neumann) entropy associated with the entanglement of the state on $B$ with its complement,
@@
The fact that outer wedges nest along the apparent horizon, i.e., $O_{W}[\mu(\tau_{2})]\subset O_{W}[\mu(\tau_{1})]$ for $\tau_{2}>\tau_{1}$, then explains the growth of the area in entropic terms, as the outer entropy is computed with more and more data about the spacetime integrated out (see, e.g., Fig. 2(b) of Ref. Engelhardt:2017aux ). A pressing open question, of particular relevance to the case of collapsing black holes and cosmological applications of holographic screens, is to provide an analogous, explicit holographic entropy construction for timelike portions of the screen. In Sec. IV, we comment on challenges inherent to doing so. However, towards that end, understanding and characterizing where HRT surfaces can be located, in generality, relative to a holographic screen is a question of critical importance to any future attempt at providing such an entropic formulation.
@@ III Sequestering holographic screens with extremal surfaces
    Cross-sectional areas of the null congruence $N_{k}[\sigma]$ are nondecreasing towards the marginal surface $\sigma$; similarly, cross-sectional areas of $N_{k}[X]$ and $N_{\ell}[X]$ are nondecreasing towards the HRT surface $X$.
@@ III.1.1 Causal future $I^{+}[X]$
Let us first focus our attention on screens with early timelike flows. In the causal future of $X$, these are screens with a timelike segment close to future infinity. By the area law for holographic screens Bousso:2015mqa ; Bousso:2015qqa , timelike segments of future holographic screens may either remain timelike and past-directed, or turn outwards into a spacelike segment; turning spacelike inwards violates monotonicity of its cross-sectional area. Since we already know that holographic screens cannot cross $N_{k}[X]$ and $N_{\ell}[X]$, if a timelike segment coming in from future infinity is to remain timelike, it can only end on or go through $X$. As shown in Fig. 4, however, it is easy to see that a forbidden loop arises in this case as well, so the only allowed screens with early timelike flows are those that have late spacelike flows that asymptote to $N_{+k}[X]$. On the other hand, by sequestration, screens with early spacelike flows in the future of $X$ must asymptote, backwards along the flow, to the null congruence $N_{+\ell}[X]$. In a similar fashion, spacelike segments of future holographic screens may either remain spacelike toward the exterior, or turn timelike towards the past, as turning timelike towards the future violates the area law. Since such screens must again not cross $N_{k}[X]$ or $N_{\ell}[X]$, or pass through $X$, they must also have late spacelike flows asymptoting to $N_{+k}[X]$ just as for the timelike case. Note that while consistency with the rules above constrain the early and late flow behavior of $H$ to the future of $X$, the screen itself is allowed to have arbitrarily many intermediate transitions between spacelike and timelike regions of $H$, as long as monotonicity as stipulated by the area law is respected.
@@ III.1.2 Outer wedge $O_{W}[X]$
In the outer wedge, by sequestration, screens that have early timelike flows must asymptote to $N_{+k}[X]$ toward the future. As we evolve along the flow, the screen can continue in a timelike fashion and asymptote to $N_{-\ell}[X]$ toward the past so as to not cross null congruences of $X$, or turn spacelike towards the outer boundary $\mathcal{B}$ in accordance with the area law. There is no penalty for the screen to cross the Cauchy surface $\Sigma$ defining $X$ as this does not generate forbidden loops, and so late spacelike flows can be either to the future or past of $\Sigma$. On the other hand, screens with early spacelike flows must begin at the HRT surface $X$ since sequestration dictates that they cannot come from another wedge. These early spacelike segments can have the same late flow behaviour as early timelike segments, i.e., they can have late spacelike flows to the future or past of $\Sigma$, or turn timelike to the past and asymptote to $N_{-\ell}[X]$. Note that screens in the outer wedge all have to end at the boundary $\mathcal{B}$.

### 07 Feb 2024 v2

@@ I Introduction
The past quarter-century of the anti–de Sitter/conformal field theory (AdS/CFT) correspondence Maldacena:1997re ; Gubser:1998bc ; Witten:1998qj ; Aharony:1999ti has ushered in a revolution of our understanding of gravitation and spacetime. Rather than being fundamental concepts, in AdS/CFT both quantum gravity and the bulk dimension of spacetime itself can be seen as emergent phenomena Susskind:1994vu ; Bousso:2002ju ; Faulkner:2013ica , arising as effective descriptions encoded via complicated and subtle details of the dynamics in the nongravitational world of the boundary. This encoding has been understood as a form of quantum error correction Almheiri:2014lwa ; Pastawski:2015qua , with the geometry encoding entanglement—specifically, the fine grained (von Neumann) entropy—through areas of surfaces, or more general quantities in a gravitational effective field theory Dong:2013qoa . [^1] See Refs. Dong:2023bax ; Gesteau:2023hbq for recent analyses of the relationship between renormalization and holographic entropy in effective theories. The Ryu-Takayanagi (RT) formula Ryu:2006bv , and more generally the Hubeny-Rangamani-Takayanagi (HRT) formalism Hubeny:2007xt for dynamical spacetimes, give the prescriptions for finding the appropriate bulk surfaces whose areas compute the entropy corresponding to a given boundary region, generalizing the celebrated Bekenstein-Hawking entropy of black holes Bekenstein:1972tm ; Bekenstein:1973ur ; Hawking:1976de .
@@
Despite the success of these famous results and many others, significant puzzles remain. How does spacetime emerge when it is not anti–de Sitter? That is, what is the analogue of the boundary theory for general spacetimes, and where is it located? As a starting point, one can look to history, where Hawking’s area theorem for black holes arguably was a critical seed leading eventually to our current holographic understanding of AdS/CFT. Happily, new area theorems for general spacetimes have been found by Bousso and Engelhardt Bousso:2015mqa ; Bousso:2015qqa , in the form of monotonic area increase along a special hypersurface called a holographic screen. The holographic screen has been conjectured to be the appropriate analogue of the AdS boundary for a holographic description of general spacetimes Bousso:1999cb ; Bousso:2015mqa ; Bousso:2015qqa ; Nomura:2017npr . Within AdS/CFT, an entropic interpretation of the outermost spacelike portion of a holographic screen, that is, an apparent horizon, was provided by Refs. Engelhardt:2017aux ; Engelhardt:2018kcs . There it was shown that the area of a slice of the screen is equal to the outer entropy: the area of the maximal HRT surface, in units of $4G\hbar$, subject to holding the geometry in the outer wedge—the exterior of the slice of the screen—fixed and marginalizing over all possible completions of the spacetime. In this sense, at least the apparent horizon can be viewed as a coarse grained holographic measure of entanglement entropy. Moreover, the outer entropy can be viewed as computing a quasilocal energy for general (i.e., nonmarginal) surfaces Bousso:2018fou ; Wang:2020vxc .
@@
It is this latter question that we answer in this paper. Specifically, we show that holographic screens are sequestered relative to the HRT surface: given an HRT surface, a holographic screen is contained entirely within either its interior, exterior, past, or future, being forever forbidden from crossing the null congruences launched from the HRT surface. There are special cases in which, under certain conditions, the screen can touch—but not pass through—a null congruence from the HRT surface. Specifically, while such instantaneous intersections are strictly forbidden for spherically symmetric geometries, in nonspherical cases they cannot be ruled out if either (i) some generators of the congruence do not intersect the screen (i.e., the intersection is incomplete), or (ii) the intersection is a surface of zero Euler characteristic. Possibility (i) is fairly generic, and indeed a complete intersection can always be infinitesimally deformed to an incomplete one. However, possibility (ii) is fairly exotic, as it requires that the HRT surface be, e.g., a topological torus, Klein bottle, or odd-dimensional sphere.
This paper is organized as follows. After providing a self-contained primer of the geometric background and definitions in Sec. II, we prove our main sequestration result, Theorem III.6, in Sec. III. We enumerate the possibilities for exceptional cases where the screen and null congruences can touch in detail in Sec. III.3. We comment on obstructions to straightforward application of a coarse grained interpretation of timelike holographic screens in Sec. IV. We conclude in Sec. V, placing our result in context and commenting on connections to non-HRT extremal surfaces (e.g., “minimax” surfaces) and semiclassical quantum corrections.
@@ II.1 Topological definitions
Throughout this paper, we will be concerned with geometric objects constructed from light rays: causally defined regions, expansion of bundles of light rays, etc. We begin, for completeness, by reviewing some terminology, following Refs. Engelhardt:2018kcs ; Nomura:2018aus ; Bousso:2018fou . For a set of points $S\subset{\cal M}$, we define the chronological and causal future of $S$ as $I^{+}[S]$ and $J^{+}[S]$, denoting the sets of all points in ${\cal M}$ connected with any point in $S$ by timelike or null paths, respectively. The chronological and causal pasts $I^{-}[S]$ and $J^{-}[S]$ are defined analogously. We will assume that ${\cal M}$ is globally hyperbolic: it is free of closed causal curves and, for all points $p,q\in{\cal M}$, $J^{+}(p)\cap J^{-}(q)$ is compact (or, in the asymptotically-AdS case, when supplemented with boundary conditions as discussed in Ref. Avis:1977yn ). We define the future (past) domain of dependence of $S$, $D^{\pm}[S]$, as the set of all $p\in{\cal M}$ for which all past (future) inextendible causal curves through $p$ pass through $S$, and define $D[S]=D^{+}[S]\cup D^{-}[S]$. We will refer to a surface as acausal if every pair of points on the surface is spacelike-separated; the weaker condition achronal is defined analogously, but allowing null-separated points. A Cauchy slice $\Sigma$ is defined as an achronal codimension-one hypersurface for which $D[\Sigma]={\cal M}$. We will at times refer to an achronal slice of the boundary ${\cal B}=\partial{\cal M}$ (i.e., ${\cal B}\cap\Sigma=\partial\Sigma$ for a Cauchy slice $\Sigma$) as itself a boundary $B$, trusting that the difference between $B$ and ${\cal B}$ will be clear in context. From a closed, codimension-two, achronal surface $\nu\subset{\cal M}$, we can define four orthogonal null congruences—families of null geodesics hereafter referred to as light sheets—associated with the null rays launched either outward or inward, past- or future-directed, from $\nu$, spatially orthogonal to $\nu$ at the launch point. Let us define the null vector on one side (e.g., “outward”) as $k$ and on the other side (“inward”) as $\ell$, with $+k$ and $+\ell$ being future-directed, and write the four congruences as $N_{\pm k}[\nu]$ and $N_{\pm\ell}[\nu]$. When convenient, we will also denote the union of future- and past-directed null congruences by $N_{k}[\nu]=N_{+k}[\nu]\cup N_{-k}[\nu]$ and $N_{\ell}[\nu]=N_{+\ell}[\nu]\cup N_{-\ell}[\nu]$. Geodesics can exit the light sheets if Wald ; Penrose and only if Akers:2017nrr they encounter either a caustic or an intersection with a distinct null geodesic. For $\nu$ Cauchy-splitting—that is, dividing a Cauchy surface $\Sigma\supset\nu$ into two sides $\Sigma^{\pm}$—we can write
@@
where $\dot{}$ denotes the boundary of a set. See Fig. 1. We will choose $\Sigma^{+}[\nu]$ to be on the side of $\nu$ toward which $k$ points and $\Sigma^{-}[\nu]$ to be on the side toward which $\ell$ points. We define the outer wedge $O_{W}[\nu]=D[\Sigma^{+}[\nu]]$ and inner wedge $I_{W}[\nu]=D[\Sigma^{-}[\nu]]$.
@@ II.2 Dynamical definitions
Having concluded the topological preliminaries, we now turn to dynamics. We will define the 2+2 formalism for general relativity, in which the Einstein equations governing the evolution of spacetime can be recast in terms of these light-centric objects. This formulation is known as the characteristic initial data formalism for general relativity Rendall:2000 ; Brady:1995na ; ChoquetBruhat:2010ih ; Luk:2011vf ; Chrusciel:2012ap ; Chrusciel:2012xf ; Chrusciel:2014lha , and one can view it as akin to the more familiar 3+1 Arnowitt-Deser-Misner formalism, but with spatial Cauchy data replaced with Cauchy data on a null surface. Given a Cauchy-splitting surface $\nu$, let us focus on the light sheet along the $k$-direction, $N_{k}[\nu]$. We will define $k$ to be affinely parametrized as it is parallel transported along itself, and $\ell$ to be parallel transported along $k$ but continually rescaled so that $k\cdot\ell=-1$. We define the induced metric as $q_{\mu\nu}=g_{\mu\nu}+k_{\mu}\ell_{\nu}+k_{\nu}\ell_{\mu}$, in terms of which we have the null extrinsic curvature,
@@ II.3 Screen-related definitions
We define a Hubeny-Rangamani-Takayanagi (HRT) surface as an extremal surface $X$ homologous to the boundary $B$, such that there exists a Cauchy slice $\Sigma$ of $D[O_{W}[X]]$ on which $X$ is minimal. As such surfaces can be identified via a maximin optimization Wall:2012uf , we will sometimes refer to HRT surfaces as “being maximin.” Though HRT surfaces can be defined for subregions of the boundary, throughout, we will be interested in taking $B$ to be a complete connected component. One can show that HRT surfaces are minimar Engelhardt:2018kcs . HRT surfaces are of particular interest to holography, since the area of the HRT surface, divided by $4G\hbar$, corresponds to the fine grained (von Neumann) entropy associated with the entanglement of the state on $B$ with its complement,
@@
The fact that outer wedges nest along the apparent horizon, i.e., $O_{W}[\mu(\tau_{2})]\subset O_{W}[\mu(\tau_{1})]$ for $\tau_{2}>\tau_{1}$, then explains the growth of the area in entropic terms, as the outer entropy is computed with more and more data about the spacetime integrated out (see, e.g., Fig. 2(b) of Ref. Engelhardt:2017aux ). A pressing open question, of particular relevance to the case of collapsing black holes and cosmological applications of holographic screens, is to provide an analogous, explicit holographic entropy construction for timelike portions of the screen. In Sec. IV, we comment on challenges inherent to doing so. However, toward that end, understanding and characterizing where HRT surfaces can be located, in generality, relative to a holographic screen is a question of critical importance to any future attempt at providing such an entropic formulation.
@@ III Sequestering holographic screens with extremal surfaces
    Cross-sectional areas of the null congruence $N_{k}[\sigma]$ are nondecreasing toward the marginal surface $\sigma$; similarly, cross-sectional areas of $N_{k}[X]$ and $N_{\ell}[X]$ are nondecreasing toward the HRT surface $X$.
@@ III.1.1 Causal future $I^{+}[X]$
Let us first focus our attention on screens with early timelike flows. In the causal future of $X$, these are screens with a timelike segment close to future infinity. By the area law for holographic screens Bousso:2015mqa ; Bousso:2015qqa , timelike segments of future holographic screens may either remain timelike and past-directed, or turn outward into a spacelike segment; turning spacelike inward violates monotonicity of its cross-sectional area. Since we already know that holographic screens cannot cross $N_{k}[X]$ and $N_{\ell}[X]$, if a timelike segment coming in from future infinity is to remain timelike, it can only end on or go through $X$. As shown in Fig. 5, however, it is easy to see that a forbidden loop arises in this case as well, so the only allowed screens with early timelike flows are those that have late spacelike flows that asymptote to $N_{+k}[X]$. On the other hand, by sequestration, screens with early spacelike flows in the future of $X$ must asymptote, backward along the flow, to the null congruence $N_{+\ell}[X]$. In a similar fashion, spacelike segments of future holographic screens may either remain spacelike toward the exterior, or turn timelike toward the past, as turning timelike toward the future violates the area law. Since such screens must again not cross $N_{k}[X]$ or $N_{\ell}[X]$, or pass through $X$, they must also have late spacelike flows asymptoting to $N_{+k}[X]$ just as for the timelike case. Note that while consistency with the rules above constrain the early and late flow behavior of $H$ to the future of $X$, the screen itself is allowed to have arbitrarily many intermediate transitions between spacelike and timelike regions of $H$, as long as monotonicity as stipulated by the area law is respected.
@@ III.1.2 Outer wedge $O_{W}[X]$
In the outer wedge, by sequestration, screens that have early timelike flows must asymptote to $N_{+k}[X]$ toward the future. As we evolve along the flow, the screen can continue in a timelike fashion and asymptote to $N_{-\ell}[X]$ toward the past so as to not cross null congruences of $X$, or turn spacelike toward the outer boundary $\mathcal{B}$ in accordance with the area law. There is no penalty for the screen to cross the Cauchy surface $\Sigma$ defining $X$ as this does not generate forbidden loops, and so late spacelike flows can be either to the future or past of $\Sigma$. On the other hand, screens with early spacelike flows must begin at the HRT surface $X$ since sequestration dictates that they cannot come from another wedge. These early spacelike segments can have the same late flow behavior as early timelike segments, i.e., they can have late spacelike flows to the future or past of $\Sigma$, or turn timelike to the past and asymptote to $N_{-\ell}[X]$. Note that screens in the outer wedge all have to end at the boundary $\mathcal{B}$.